Multiple-Level Break-Even Analysis
Nielsen Associates provides marketing services for a number of small manufacturing firms. Nielsen receives a commission of 10 percent of sales. Operating costs are as follows:

Unit-level costs $0.02 per sales dollar
Sales-level costs $200 per sales order
Customer-level costs $800 per customer per year
Facility-level costs $60,000 per year

(a) Determine the minimum order size in sales dollars for Nielsen to break even on an order.
$Answer



(b) Assuming an average customer places five orders per year, determine the minimum annual sales required to break even on a customer.
$Answer



(c) What is the average order size in (b)?
$Answer



(d) Assuming Nielsen currently serves 100 customers, with each placing an average of five orders per year, determine the minimum annual sales required to break even.
$Answer



(e) What is the average order size in (d)?
$Answer



(f) Explain the differences in the answers to (a), (c), and (e).

In multiple customer firms the break-even point decreases as the number of customers increases.

Even if individual orders have a positive contribution, some customers may be unprofitable.

In the long-run the most important costs are facility level costs.

The most important costs to cover are unit level costs.
Answer #1
(a) Determine the minimum order size in sales dollars for Nielsen to break even on an order.
Lets assume X is the minimum order size
X = $200 + $0.02X
0.08 X = 200 2500.00
Minimum Order Size = 2500.00
(b) Assuming an average customer places five orders per year, determine the minimum annual sales required to break even on a customer.
Lets assume that X is the minimum annual sales to a customer
X = 800 + 5 * 200 + 0.02 * X
0.08X = 1,800 22500.00
Minimum annual sales required to break even on a customer 22500.00
(c) What is the average order size in (b)?
Average order size = Minimum annual sales required to break even on a customer /Number of orders placed by customer = 22500/5 4500.00
(d) Assuming Nielsen currently serves 100 customers, with each placing an average of five orders per year, determine the minimum annual sales required to break even.
Lets assume Y is Minimum annual sales required to break even
Y = 60,000 + 800 x 100 + 200 x  100 x  5 + 0.02 x* Y
0.08Y = 240,000 $ 3,000,000.00
Minimum annual sales required to break even = $ 3,000,000.00
(e) What is the average order size in (d)?
Avrage order size  = Average annual sales/ Number of ordes
Avrage order size  = $3,000,000/(100 x 5) 6000.00
(f) Explain the differences in the answers to (a), (c), and (e).
In multiple customer firms the break-even point decreases as the number of customers increases.
Even if individual orders have a positive contribution, some customers may be unprofitable.
In the long-run the most important costs are facility level costs.
The most important costs to cover are unit level costs. correct